OpenStudy (anonymous):

what is -7 1/4+4 2/3?

OpenStudy (anonymous):

is it\[-7 \frac{ 1 }{ 4 }+4 \frac{ 2 }{ 3 }\]?

OpenStudy (anonymous):

\[-\frac{ 31 }{ 12 }\]

OpenStudy (anonymous):

i got it as -2 7/12

OpenStudy (anonymous):

well it is technically the same thing if you say 4 2/3 - 7 1/4

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

did you find a common denominator?

OpenStudy (anonymous):

yes and it is 12

OpenStudy (anonymous):

right